sqrt (x-1)= 4 sqrt (x+1)
Answer by fractalier(6550)

From
sqrt (x-1)= 4 sqrt (x+1)
we square both sides and get
x - 1 = 16(x + 1)
x - 1 = 16x + 16
Rearranging we get
-17 = 15x
x = -17/15
But notice this gives unreal numbers, as you cannot have real roots of negative numbers...therefore the answer is
No Solution (in the real numbers)